# Structure-Preserving Neural Ordinary Differential Equations for Stiff Systems

### v3

@@ 1 Introduction
Data-driven reduced order models (ROMs) of dynamical systems arising from high-resolution simulations or experiments can be used as efficient forward-model surrogates. The learned surrogates can be deployed in the traditionally expensive tasks of optimization and parameter inference. However, conventional data-driven ROMs for dynamical systems generally struggle with accurately approximating stiff dynamical systems [1] and maintaining long-term stability [2, 3]. In this work, we address both of these challenges in the context of neural ordinary differential equations (NODEs) [4], which are a class of models that approximate dynamical systems continously as ODEs using neural networks. The two ingredients that determine the effectiveness of NODE-based models are the neural network architecture and the numerical integrator. For the former we proposed a structure-preserving NODE, which integrates with a linear and nonlinear split and several small components that preserve key mathematical structures strongly. For the latter we incorporate NODE with advanced time stepping through exponential integrators. The use of an exponential integrator improves the performance of explicit and semi-implicit integrators that are commonly used in packages such as diffrax [5] and torchode [6].
@@
In [2], motivated by the inertial manifold theorem, we developed a special NODE approach that embeds a linear and nonlinear partition of the right-hand side (RHS) being approximated by two neural networks. This partitioning method provides better short-time tracking, prediction of the energy spectrum, and robustness to noisy initial conditions than standard NODEs. The recent work [7, 8] extended the linear and nonlinear partition to a semi-implicit NODE. In the forward pass of their training they solve a partitioned ODE using implicit-explicit Runge-Kutta (IMEX-RK) methods. The linear and nonlinear partition allows them to avoid solving a fully nonlinear system. The backward pass is handled through a discrete adjoint approach, resulting into a linear system to be solved. In general, the matrices from these linear systems are dense, requiring a direct solve which can be computationally expensive and have poor scalability. Our NODE uses an explicit approach based on exponential integrators, which do not require direct solvers but enjoy stability properties comparable to implicit approaches. To the best of our knowledge, exponential integration schemes have first been incorporated into NODEs in recent work [9] The authors show that exponential integrators can outperform implicit schemes for low-dimensional ODEs. However, this work shows the straightforward implementation of exponential integrators for high-dimensional systems to be too computationally expensive for practical purposes. Our work relies on several key components to overcome the challenge of training when applying such an exponential integrator with NODE to high-dimensional problems.
@@
The contributions of our work include both structure-preserving enhancements and improvements in time stepping and training. In our approach, a coordinate transformation is learned via an autoencoder in which the dynamics are described by a NODE with a linear and nonlinear partitioning. The linear part of the system is parameterized using a novel Hurwitz matrix decomposition based on [10]. This promotes long-term stability through bounding the real parts of the linear operator eigenvalues from above by a specified shift, $\mu\in\mathbb{R}$. The nonlinear operator is represented using either a bilinear form or a Lipshitz controlled neural network based on the innovations of our previous work [3]. The bilinear form has the advantage of clearly separating the linear and nonlinear operators while a neural network still contains a nonzero linear contribution in it’s Taylor expansion. In either case, we show that a local Lipschitz bound on the nonlinear operator can provably guarantee stable long-term predictions for certain dynamical systems. Our other major contribution involves using an exponential integrator for the time integration. We propose a matrix-free method for advancing the dynamics based on the algorithm developed in [11]. In our numerical examples, we show that the combination of structure-preserving networks and exponential integrators enables us to robustly learn the Robertson system, which is a stiff system bridging time scales over multiple orders of magnitude. For cases with high-dimensional data we use an autoencoder to significantly reduce the dimensionality and aid in our training. We demonstrate using the Kuromoto-Sivashinsky equations that evolving dynamics on a low-dimensional latent space while preserving the linear and nonlinear partition allows a significant reduction in training time while retaining the stability of the model. We note that latent dynamics discovery has seen success in other scientific machine learning approaches (e.g. [12, 13]).
@@
Training and deploying neural ODEs in practical problems is often complicated by stiffness in the dynamics, which occurs when some or all of the solution components undergo rapid changes relative to transient behavior. Stiff behavior forces explicit solvers to take prohibitively small timesteps. A number of recent works have utilized ML to address the challenges of stiff systems. In [14] the stiffness issue is resolved by introducing a suitable reparameterization in time. The map produces a nonstiff surrogate that can be cheaply solved with an explicit method. In [13] the authors utilize a physics-assisted autoencoder to learn a low-dimensonal latent space of a collisional radiative model. In [1] the authors utilize equation scaling to mitigate the stiffness problem. The method proposed in this work is capable of resolving the stiff dynamics by learning a suitable coordinate transformation into a new system composed of a linear operator capturing the stiffness of the system and a nonlinear operator which can be easily integrated using a standard first-order exponential integrator. This capability is highlighted in Roberton’s chemical reaction problem in Section 5.2, in which conventional exponential integrators fail to eliminate the stiff time step restriction resulting from the original form of the equations.
@@
The rest of the paper is organized as follows. Section 2 presents the details behind the parameterizations of the linear and nonlinear parts of NODE along with implementation details. Section 3 gives a result about the Lyapunov stability and provides an error bound for the NODE. Section 4 provides a brief introduction to exponential integrators. The details related to training and their results are given in Section 5. Conclusive remarks are given in 6.
@@ 2.1 Hurwitz Matrix Parameterization
This theorem and its proof can be found in [10].
@@ 2.2 Bilinear Form Neural Network
One option to approximate the nonlinear part, $g$, is to use a bilinear form. By construction, the bilinear form has zero contribution to the linear operator since it is built with quadratic forms. This property is not guaranteed when using a vanilla feed-forward network for $g$. This creates a strict separation of the linear and nonlinear dynamics that allows recovery of the linear operator $A$ in an equation learning setting. We utilize the low rank-bilinear form network as defined in [3]. It is defined by
@@
[…]
@@ 2.3 Lipschitz Controlled Neural Network
First, we show each layer $l_{i}$ is $L^{1/p}$-Lipschitz. If $z_{1}$ and $z_{2}$ are inputs for $l_{i}$, then
@@
Our previous work [3] introduced a bi-Lipschitz affine transformation (bLAT) network, achieving Lipschitz control using a parameterization based on the singular value decomposition (SVD). The parameteization presented here is more efficient since the SVD is more expensive to evaluate than the $1$ and $\infty$ norms. However, unlike the above, the bLAT network is bi-Lipschitz.
@@ 2.4 Autoencoder
Inspired by the seminal work [15] for complex dynamical systems involving high-dimensional data, an autoencoder produces a low dimensional representation of the dynamics which can be integrated efficiently. For high-dimensional problems like the Kuramoto-Sivashinsky equation, we find it is necessary to include the dimension reduction in structure-preserving NODE to avoid training directly on a high-dimensional space, which is more challenging and expensive. For low-dimensional problems such as the Robertson problem in Section 5.2, the autoencoder is additionally used to learn a transformation into coordinates that yields a system of the same dimension but with stiffness coming from the linear part. Instead of simplifying our model by reducing the dimension, the system is transformed such that our exponential integrator is able to integrate utilizing large timesteps.
We consider some dynamical system for $U(t)\in\mathbb{R}^{m}$ and assume that it can be transformed into a variable $u\in\mathbb{R}^{n}$, $n<m$ where the dynamics are described by a system of the form (1). Two feed-forward neural networks are trained $f_{E}(U;\alpha)$ and $f_{D}(u;\beta)$, which represent the encoder and decoder and have network parameters $\alpha$ and $\beta$, respectively. The network $f_{E}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ takes in the initial condition in physical space and network parameters and outputs the initial condition in the latent space $\mathbb{R}^{n}$ with $n\leq m$. Once the initial condition is projected onto the latent space we use a time integrator to evolve the dynamics forward in time. By evolving the dynamics in the low-dimensional latent space $\mathbb{R}^{n}$, efficiency is gained in the numerical integration step. In mathematical symbols we have $f_{E}(U_{0})=u_{0}$ then we evolve $u_{0}$ several timesteps (using the exponential integrator in this work) to obtain $\{u_{1},u_{2},\dots,u_{k}\}$. After this, these are each transformed back into the full space data using the decoder $U_{i}=f_{D}(u_{i})$, which are then used to compute the loss term. A schematic of this process is given in Figure 1.
Figure 1: Schematic of the autoencoder for our structure preserving NODE. The dark blue block represents the initial data which is encoded using $f_{E}$ into the light blue block representing the initial condition in the latent space. The green and purple blocks represent the neural networks for the linear and nonlinear operators, respectively. These neural networks are used to push the dynamics forward using the exponential integrator defined in (17). The data in the latent space is then decoded back into physical space using $f_{D}$. Decoded data is represented by the pink blocks.
@@ 3.1 Lyapunov Stability of the Network
$\frac{d}{dt}\tilde{u}=A\tilde{u}+\tilde{g}_{NN}(\tilde{u})$ (9)
where $\|\tilde{g}_{NN}\|\leq L\|\tilde{u}\|$ near $\tilde{u}=0$. Therefore, without loss of generality, we prove the theorem using the notation of the original system assuming $u_{e}=0$ and $\|g\|\leq L\|u\|$. It suffices to show the existence of a Lyapunov function, $V(u)$, such that $V(u)\geq 0,V(u)=0$ if and only if $u=0$, and $\frac{d}{dt}V(u)\leq 0$ for all $t>0$. Let $A_{L}=TJT^{-1}$ be the Jordan decomposition for $A_{L}$. According to [16], for any $\epsilon>0$ it is possible to write
$J=T^{-1}A_{L}T=\rm{diag}\left\{\begin{pmatrix}\lambda_{1}&\epsilon&\\ &\ddots&\epsilon\\ &&\lambda_{1}\end{pmatrix},\begin{pmatrix}\lambda_{2}&\epsilon&\\ &\ddots&\epsilon\\ &&\lambda_{2}\end{pmatrix},\dots\right\},$ (10)
@@ 4 Exponential Integrators
[…]
@@ A $p$-th order Taylor series expansion of $g$ about $t_{0}$ gives
[…]
@@
The integrator used in this study is exponential time differencing (ETD) of order 1 [18]. The ETD methods are based on approximating the nonlinear term by an algebraic polynomial. Note that there are many more exotic versions of exponential integrators such as exponential Rosenbrock methods and Lawson methods [17].
Evaluating the $\varphi$ functions is a nontrivial task. The computation of $\varphi_{1}$ is a well known problem in numerical analysis [19, 20, 21]. Ref. [11] shows that it is possible to avoid computing the $\varphi$ functions (14) by solving a system with a single matrix exponential of size $(n+p)\times(n+p)$, where $n$ is the dimension of the system and $p$ is the order of the approximation. The system (1) is approximated using
@@
where $e_{p}$ is a vector with a $1$ in the $p^{\rm th}$ element and 00’s elsewhere and $W\in\mathbb{R}^{n\times p}$ and $J\in\mathbb{R}^{p\times p}$ are defined to be
@@
A matrix-free algorithm developed in [11] is used to compute (17). This algorithm adapts the scaling and squaring method [22] by computing $e^{tA}b\approx(T_{m}(s^{-1}tA))^{s}b$, where $T_{m}$ is a truncated Taylor series. Algorithmically, the computation is performed using
@@
The choice of $s$ and $m$ impact both accuracy and efficiency. The number of matrix vector multiplications is $sm$ and an analysis for the optimal choice of $s$ and $m$ may be found in [11]. In practice, we choose $s$ to be proportional to the maximum time step of the training data and the norm of $A$, $s\sim\Delta t_{\rm max}\|A\|_{2}$, and we choose $m$ such that (19c) converges to some tolerance, $\tau$. Utilizing the matrix-free computation eliminates the need to compute the matrix exponential in equation (15), and can provide a faster alternative when the dimension of $A$ is large. Moreover, the computation of a matrix exponential is required for each trajectory due to the time dependency. Batching of data is improved since a matrix-free algorithm allows for vectorization. This form is used to evolve the dynamics for the structure-preserving NODE in Section 5. Note that higher-order versions of this scheme require the derivatives of the solution.

### v4

@@ 1 Introduction
Data-driven reduced order models (ROMs) of dynamical systems arising from high-resolution simulations or experiments can be used as efficient forward-model surrogates. The learned surrogates can be deployed in the traditionally expensive tasks of optimization and parameter inference. However, conventional data-driven ROMs for dynamical systems generally struggle with accurately approximating stiff dynamical systems kim2021stiff and maintaining long-term stability linot2023stabilized, serino2024intelligent. In this work, we address both of these challenges in the context of neural ordinary differential equations (NODEs) chen2018neural, which are a class of models that approximate dynamical systems continuously as ODEs using neural networks. The two ingredients that determine the effectiveness of NODE-based models are the neural network architecture and the numerical integrator. For the former we proposed a structure-preserving NODE, which integrates with a linear and nonlinear split and several small components that preserve key mathematical structures strongly. For the latter we incorporate NODE with advanced time stepping through exponential integrators. The use of an exponential integrator improves the performance of explicit and semi-implicit integrators that are commonly used in packages such as diffrax kidger2021on and torchode lienen2022torchode.
@@
In linot2023stabilized, motivated by the inertial manifold theorem, we developed a special NODE approach that embeds a linear and nonlinear partition of the right-hand side (RHS) being approximated by two neural networks. This partitioning method provides better short-time tracking, prediction of the energy spectrum, and robustness to noisy initial conditions than standard NODEs. The recent work zhang2023semi, zhang2024semi extended the linear and nonlinear partition to a semi-implicit NODE. In the forward pass of their training they solve a partitioned ODE using implicit-explicit Runge-Kutta (IMEX-RK) methods. The linear and nonlinear partition allows them to avoid solving a fully nonlinear system. The backward pass is handled through a discrete adjoint approach, resulting into a linear system to be solved. In general, the matrices from these linear systems are dense, requiring a direct solve which can be computationally expensive and have poor scalability. Our NODE uses an explicit approach based on exponential integrators, which do not require direct solvers but enjoy stability properties comparable to implicit approaches. To the best of our knowledge, exponential integration schemes have first been incorporated into NODEs in recent work fronk2024training The authors show that exponential integrators can outperform implicit schemes for low-dimensional ODEs. However, this work shows the straightforward implementation of exponential integrators for high-dimensional systems to be too computationally expensive for practical purposes. Our work relies on several key components to overcome the challenge of training when applying such an exponential integrator with NODE to high-dimensional problems.
Figure 1: Schematic of the autoencoder for our structure preserving NODE. The dark grey block represents the initial data which is encoded using $f_{E}$ represented by the left light grey block. The latent space dynamics are evolved in the light blue block. The green and purple blocks represent the neural networks for the linear and nonlinear operators, respectively. These neural networks are used to push the dynamics forward using the exponential integrator defined in (17) represented by the light orange colored blocks. The data in the latent space is then decoded back into physical space using $f_{D}$. Decoded data is represented by the right light grey block. The solution, $U_{t}$ is represented by the right most block.
The contributions of our work include both structure-preserving enhancements and improvements in time stepping and training. In our approach, a coordinate transformation is learned via an autoencoder in which the dynamics are described by a NODE with a linear and nonlinear partitioning. Our approach is summarized by the schematic in Figure 1. In related work, augmented Neural ODEs dupont2019 expand the hidden state space to enhance expressivity and improve numerical stability. Likewise, the structure-preserving formulation introduced here implicitly augments the state space through a linear-nonlinear partition, achieving similar advantages for modeling stiff systems.
@@
The linear part of the system is parameterized using a novel Hurwitz matrix decomposition based on duan1998note. This promotes long-term stability through bounding the real parts of the linear operator eigenvalues from above by a specified shift, $\mu\in\mathbb{R}$. The nonlinear operator is represented using either a bilinear form or a Lipshitz controlled neural network based on the innovations of our previous work serino2024intelligent. The bilinear form has the advantage of clearly separating the linear and nonlinear operators while a neural network still contains a nonzero linear contribution in it’s Taylor expansion. In either case, we show that a local Lipschitz bound on the nonlinear operator can provably guarantee stable long-term predictions for certain dynamical systems. Lyapunov-based neural networks lyapunov24 offer a related perspective, as they enforce stability through explicit Lyapunov-function parameterizations that guarantee monotonic energy decay along system trajectories. While sharing the goal of stability preservation, these methods are rooted in control-oriented formulations, whereas our approach achieves Lyapunov stability implicitly through Hurwitz spectral constraints on the linear operator and Lipschitz control on the nonlinear term, tailored specifically for stiff dynamical systems. The other major contribution of our work involves using an exponential integrator for the time integration. We propose a matrix-free method for advancing the dynamics based on the algorithm developed in al2011computing. In our numerical examples, we show that the combination of structure-preserving networks and exponential integrators enables us to robustly learn the Robertson system, which is a stiff system bridging time scales over multiple orders of magnitude. For cases with high-dimensional data we use an autoencoder to significantly reduce the dimensionality and aid in our training. We demonstrate using the Kuromoto-Sivashinsky equations that evolving dynamics on a low-dimensional latent space while preserving the linear and nonlinear partition allows a significant reduction in training time while retaining the stability of the model. We note that latent dynamics discovery has seen success in other scientific machine learning approaches (e.g. fries2022lasdi, xie2024latent).
@@
Training and deploying neural ODEs in practical problems is often complicated by stiffness in the dynamics, which occurs when some or all of the solution components undergo rapid changes relative to transient behavior. Stiff behavior forces explicit solvers to take prohibitively small timesteps. A number of recent works have utilized ML to address the challenges of stiff systems. In caldana2024neural the stiffness issue is resolved by introducing a suitable reparameterization in time. The map produces a nonstiff surrogate that can be cheaply solved with an explicit method. In xie2024latent the authors utilize a physics-assisted autoencoder to learn a low-dimensonal latent space of a collisional radiative model. In kim2021stiff the authors utilize equation scaling to mitigate the stiffness problem. The method proposed in this work is capable of resolving the stiff dynamics by learning a suitable coordinate transformation into a new system composed of a linear operator capturing the stiffness of the system and a nonlinear operator which can be easily integrated using a standard first-order exponential integrator. This capability is highlighted in Roberton’s chemical reaction problem in Section 5.2, in which conventional exponential integrators fail to eliminate the stiff time step restriction resulting from the original form of the equations.
@@
The rest of the paper is organized as follows. Section 2 presents the details behind the parameterizations of the linear and nonlinear parts of NODE along with implementation details. Section 3 gives a result about the Lyapunov stability and provides an error bound for the NODE. Section 4 provides a brief introduction to exponential integrators. The details related to training and their results are given in Section 5. Conclusive remarks are given in 7.
@@ 2.1 Hurwitz Matrix Parameterization
This theorem and its proof can be found in duan1998note.
@@ 2.2 Bilinear Form Neural Network
One option to approximate the nonlinear part, $g$, is to use a bilinear form. By construction, the bilinear form has zero contribution to the linear operator since it is built with quadratic forms. This property is not guaranteed when using a vanilla feed-forward network for $g$. This creates a strict separation of the linear and nonlinear dynamics that allows recovery of the linear operator $A$ in an equation learning setting. We utilize the low rank-bilinear form network as defined in serino2024intelligent. It is defined by
@@
where $u\in\mathbb{R}^{n},\ C^{(r)}\in\mathbb{R}^{n\times n},\ D^{(r)}\in\mathbb{R}^{n\times n},$ $b\in\mathbb{R}^{n\times 1}$, $\odot$ represents the Hadamard product (element-wise multiplication), and $R,\ 1\leq R\leq n$ represents the rank of approximation which is chosen much smaller than $n$ during the training.
@@ 2.3 Lipschitz Controlled Neural Network
First, we show each layer $l_{i}$ is $L^{1/p}$-Lipschitz. Let $z_{1}$ and $z_{2}$ be inputs for the layer $l_{i}$. Since $\sigma$ is a 1-Lipschitz activation function,
@@
Our previous work serino2024intelligent introduced a bi-Lipschitz affine transformation (bLAT) network, achieving Lipschitz control using a parameterization based on the singular value decomposition (SVD). The parameteization presented here is more efficient since the SVD is more expensive to evaluate than the $1$ and $\infty$ norms. However, unlike the above, the bLAT network is bi-Lipschitz.
@@ 2.4 Autoencoder
Inspired by the seminal work lusch2018deep for complex dynamical systems involving high-dimensional data, an autoencoder produces a low dimensional representation of the dynamics which can be integrated efficiently. For high-dimensional problems like the Kuramoto-Sivashinsky equation, we find it is necessary to include the dimension reduction in structure-preserving NODE to avoid training directly on a high-dimensional space, which is more challenging and expensive. For low-dimensional problems such as the Robertson problem in Section 5.2, the autoencoder is additionally used to learn a transformation into coordinates that yields a system of the same dimension but with stiffness coming from the linear part. Instead of simplifying our model by reducing the dimension, the system is transformed such that our exponential integrator is able to integrate utilizing large timesteps.
We consider some dynamical system for $U(t)\in\mathbb{R}^{m}$ and assume that it can be transformed into a variable $u\in\mathbb{R}^{n}$, $n<m$ where the dynamics are described by a system of the form (1). Two feed-forward neural networks are trained $f_{E}(U;\alpha)$ and $f_{D}(u;\beta)$, which represent the encoder and decoder and have network parameters $\alpha$ and $\beta$, respectively. The network $f_{E}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ takes in the initial condition in physical space and network parameters and outputs the initial condition in the latent space $\mathbb{R}^{n}$ with $n\leq m$. Once the initial condition is projected onto the latent space we use a time integrator to evolve the dynamics forward in time. By evolving the dynamics in the low-dimensional latent space $\mathbb{R}^{n}$, efficiency is gained in the numerical integration step. In mathematical symbols we have $f_{E}(U_{0})=u_{0}$ then we evolve $u_{0}$ several timesteps (using the exponential integrator in this work) to obtain $\{u_{1},u_{2},\dots,u_{t}\}$. After this, these are each transformed back into the full space data using the decoder $U_{i}=f_{D}(u_{i})$, which are then used to compute the loss term. A schematic of this process is given in Figure 1.
@@ 3.1 Lyapunov Stability of the Network
where $\|\tilde{g}_{NN}\|\leq L\|\tilde{u}\|$ near $\tilde{u}=0$. Therefore, without loss of generality, we prove the theorem using the notation of the original system assuming $u_{e}=0$ and $\|g\|\leq L\|u\|$. It suffices to show the existence of a Lyapunov function, $V(u)$, such that $V(u)\geq 0,V(u)=0$ if and only if $u=0$, and $\frac{d}{dt}V(u)\leq 0$ for all $t>0$. Let $A_{L}=TJT^{-1}$ be the Jordan decomposition for $A_{L}$. According to harier1993, for any $\epsilon>0$ it is possible to write
$J=T^{-1}A_{L}T=\rm{diag}\left\{\begin{pmatrix}\lambda_{1}&\epsilon&\\ &\ddots&\epsilon\\ &&\lambda_{1}\end{pmatrix},\begin{pmatrix}\lambda_{2}&\epsilon&\\ &\ddots&\epsilon\\ &&\lambda_{2}\end{pmatrix}$$\dots\right\},$ (10)
where $\lambda_{1},\lambda_{2},\dots$ are the unique eigenvalues of $A_{L}$. Let $z=T^{-1}u\in\mathbb{C}^{n}$. We define $V(u)=\langle z,z\rangle=z^{*}z$ where $\langle\cdot,\cdot\rangle$ is the complex inner product, then
@@
It is important to note that the above Lyapunov stability result of the proposed network is only applicable to a neighborhood of an equilibrium. It does not extend to other attractors or chaotic systems. Nevertheless, we show later empirically that this predicts a good long-time solution for chaotic dynamics when evaluated using appropriate statistical measures. The above assumption of addressing network stability near equilibrium is motivated by the practical stiff dynamics such as chemical reaction and collisional radiative models xie2024latent, where the dynamics of interest is a transition from various initial conditions to the corresponding equilibria.
@@ 4 Exponential Integrators
The proposed NODE form (1) assumes a linear and nonlinear split where the stiffness of the system shall come primarily from the linear matrix $A$. For such a system, the exponential integrator is an ideal technique to overcome its stiffness, and therefore it is leveraged in this work in the NODE training. In this section a brief review exponential integrators is presented. For an in depth introduction see hochbruck2010exponential. Consider a system that may be written in the form
@@ A $p$-th order Taylor series expansion of $g$ about $t_{0}$ gives
$g_{k}=\left.\frac{d^{k-1}}{dt^{k-1}}g(t,u(t))\right|_{t=t_{0}},\quad\varphi_{k}(z)=\frac{1}{(k-1)!}\int\limits_{0}^{1}e^{(1-\theta)z}\theta^{k-1}\ d\theta,\quad k\geq 1.$
@@
The integrator used in this study is exponential time differencing (ETD) of order 1 minchev2005review. The ETD methods are based on approximating the nonlinear term by an algebraic polynomial. Note that there are many more exotic versions of exponential integrators such as exponential Rosenbrock methods and Lawson methods hochbruck2010exponential.
Evaluating the $\varphi$ functions is a nontrivial task. The computation of $\varphi_{1}$ is a well known problem in numerical analysis higham2002accuracy, hochbruck1998exponential, lu2003computing. Ref. al2011computing shows that it is possible to avoid computing the $\varphi$ functions (14) by solving a system with a single matrix exponential of size $(n+p)\times(n+p)$, where $n$ is the dimension of the system and $p$ is the order of the approximation. The system (1) is approximated using
@@
where $e_{p}$ is a vector with a $1$ in the $p^{\rm th}$ element and 0’s elsewhere and $W\in\mathbb{R}^{n\times p}$ and $J\in\mathbb{R}^{p\times p}$ are defined to be
@@
A matrix-free algorithm developed in al2011computing is used to compute (17). This algorithm adapts the scaling and squaring method higham2005scaling by computing $e^{tA}b\approx(T_{m}(s^{-1}tA))^{s}b$, where $T_{m}$ is a truncated Taylor series. Algorithmically, the computation is performed using
@@
The choice of $s$ and $m$ impact both accuracy and efficiency. The number of matrix vector multiplications is $sm$ and an analysis for the optimal choice of $s$ and $m$ may be found in al2011computing. In practice, we choose $s$ to be proportional to the maximum time step of the training data and the norm of $A$, $s\sim\Delta t_{\rm max}\|A\|_{2}$, and we choose $m$ such that (19c) converges to some tolerance, $\tau$. Utilizing the matrix-free computation eliminates the need to compute the matrix exponential in equation (15), and can provide a faster alternative when the dimension of $A$ is large. Moreover, the computation of a matrix exponential is required for each trajectory due to the time dependency. Batching of data is improved since a matrix-free algorithm allows for vectorization. This form is used to evolve the dynamics for the structure-preserving NODE in Section 5. Note that higher-order versions of this scheme require the derivatives of the solution.
The matrix-free exponential integrator used in this work is currently implemented as a prototype in high-level Python, and its performance therefore reflects an unoptimized version of the algorithm. Its computational overhead depends on factors such as the number of matrix-vector multiplications, the system dimension, and the convergence properties of the chosen polynomial or rational approximation, which in turn depend on the spectrum of the matrix. In practice, the matrix-free approach scales favorably for large or sparse systems where direct matrix exponentiation or repeated linear solves are computationally expensive. Future compiled or GPU-accelerated implementations have the potential to substantially reduce the overhead, and we expect that further development of this algorithm by the numerical analysis community could yield significant performance gains.